OpenStudy (anonymous):

Please HELP. MEDALS given the graph, what is the transformed parent functions equation?

OpenStudy (anonymous):

Also how do I find dilation ror stretch!

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I need to write the equation to the transformed parent function

OpenStudy (anonymous):

Any help pls

OpenStudy (noelgreco):

you still need this?

OpenStudy (anonymous):

@NoelGreco yes pls

OpenStudy (anonymous):

@amistre64 any help?

OpenStudy (amistre64):

which one are we trying for?

OpenStudy (anonymous):

first one #13 :D or both

OpenStudy (anonymous):

the first link

OpenStudy (anonymous):

i

OpenStudy (amistre64):

and you feel this is some e^x abomination eh

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my friend thinks that, but i think its a sqrt function.. idk

OpenStudy (amistre64):

i cant see the original setup thru the pencil marks .... is that bottom right a stopping point?

OpenStudy (anonymous):

i dont think so, these are parent functions.

OpenStudy (anonymous):

unless it's a sqrt, nvmd

OpenStudy (amistre64):

y = ax^2, but at x=-3, y=1 1 = 9a a = 1/9 -5,4 looks like another reference point 4 = (-5)^2/9 36 = 25 doesnt seem to fit for a sqrt function to me

OpenStudy (anonymous):

at -4, maybe that's a horizontal asymtote- hence the expotential?

OpenStudy (amistre64):

maybe, but i cant tell what that was originally thru the pencil mark

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

oh wait, ill upload another pic

OpenStudy (amistre64):

that ones better, it looks like its hooking to me, turning upward at 7,-4 so an exponential is not likely

OpenStudy (amistre64):

looks like an x^2 modification to me at the moment

OpenStudy (anonymous):

oh, then it's sqrt? but i can seem to find an equation for the reflection..

OpenStudy (amistre64):

its not a sqrt, a sqrt doesnt have a hook, it simply ends.

OpenStudy (anonymous):

oh, i see your point..i never thought it would be quadratic.

OpenStudy (amistre64):